# Less is More: DocString Compression in Code Generation

### v2

@@ 5.3. RQ3: What is the impact of compressed prompts on developers? (Human Study)
RQ1 and RQ2 rely on automated evaluation metrics to compare the performance of different methods. In RQ3, we aim to gain more insights by conducting a case study to assess the qualitative impact of DocString compression using human evaluation.
To begin with, we present a case study to visually contrast the compression effects of our tool with other methods. Fig. 3 illustrates this by showcasing a DocString from the Llama3.1 model within the MBPP dataset. We display the original DocString alongside its compressed versions, processed by four distinct methods. The original DocString led to semantically incorrect code generation by the Llama3.1 model. The Random compression method resulted in irrelevant code. Both Selective_Context and LLMLingua2 maintained the model’s output, yet failed to rectify the semantic inaccuracy. In stark contrast, our tool’s compression successfully steered the model towards semantically correct code generation.
To further evaluate the quality of the compressed DocStrings, we conducted a human study using three groups of DocStrings compressed by different methods (Selective_Context, LLMLingua2, and ShortenDoc).
@@
Figure 4. A Sample Questionnaire Used in Human Study
To conduct the human study, we recruited three evaluators with a strong background in Python programming (all Ph.D. candidates). We selected all samples from the six datasets used in the previous experiments as the evaluation subjects. For each sample, we collected the original DocString along with the three compressed DocStrings generated by three methods.
To ensure comprehensive evaluation, the samples were divided into six groups according to the dataset, and each group was evaluated by the three evaluators. Each DocString was assessed anonymously in terms of both Informativeness and Comprehensibility. The evaluation scale ranged from 0 to 4 points, where a higher score reflected higher quality. The final score for each criterion was the average of the scores from all three evaluators.
The questionnaire used for the study is illustrated in the Fig. 4. To maintain the quality of the human evaluation, the compressed DocStrings were presented in random order, ensuring that evaluators were unaware of which method was used to generate each DocString. Additionally, the evaluators were allowed to use the internet to look up unfamiliar concepts, ensuring that they had access to relevant information. To prevent evaluator fatigue and maintain focus, we limited each evaluator to assessing no more than 20 samples per half-day session.
@@
Aspect Method HumanEval CodeHarmony MBPP Subtle Creative BigCodeBench
@@
In terms of Informativeness, the ShortenDoc method consistently scores higher than the other two methods across most datasets. This suggests that ShortenDoc is adept at preserving the essential information within the DocStrings even after compression. In addition, the high Fleiss’ Kappa scores, ranging from 0.7 to 0.9, indicate a strong consensus among the evaluators regarding the Informativeness of the compressed DocStrings. This high agreement likely stems from the objective nature of this metric, where evaluators can clearly discern whether the compressed DocString retains the necessary information from the original.
In terms of Comprehensibility, ShortenDoc also shows superiority than the other two methods in Comprehensibility, indicating that the compressed DocStrings remain understandable to human readers. While the lower Fleiss’ Kappa scores for Comprehensibility, ranging from 0.1 to 0.3, suggest that there is less consistency among the evaluators in this metric compared to Informativeness. This variability could be attributed to several factors: (1) Subjectivity. Comprehensibility is inherently more subjective than Informativeness, as it depends on the evaluator’s personal ability to understand the compressed information. (2) Complexity of Docstrings. Some DocStrings, despite being compressed, may still contain complex information or jargon that is more challenging for some evaluators to grasp. (3) Diversity in Evaluator Background. The professional background and experience of the evaluators can influence their assessment of DocString comprehensibility. For instance, evaluators with specific domain knowledge might find certain terms or concepts more comprehensible.
@@ 6.1. The Impact of Method Name Quality
Table 7. Efficacy comparison within Different Method Name Quality
DataSet Method DS-1.3B DS-6.7B CQ-7.3B CG-9.4B LA-8.0B Avg.
@@
To empirically validate this hypothesis, we designed a controlled experiment. We systematically replaced high-quality method names with a generic placeholder, such as ‘foo’ (Yang et al., 2024b), across various datasets including HumanEval, CodeHarmony, MBPP, Subtle, and Creative.
@@
The outcomes of this experiment are encapsulated in Table 7. The data clearly demonstrate that method names wield a significant influence on the efficacy of DocString compression. Across the board, models exhibited a notable decline in efficacy when high-quality method names were replaced with the generic ‘foo’. This finding suggests that a portion of the information compressed by ShortenDoc is redundant with the information already present in the method name.
Conversely, when low-quality method names were enhanced, we observed a consistent uptick in model efficacy. This improvement underscores the notion that enriching the quality of method names not only aids in the comprehension of the code but also bolsters the robustness of ShortenDoc’s compression capabilities.
@@ 6.2. The Impact of DocString Style
Experiments with Newline and Tab Characters. In Section 4, during the preprocessing phase, we made an assumption that the removal of newline characters and tabs would not substantially affect the model’s code generation capabilities. This assumption was based on the idea that these characters, while useful for formatting, might not contribute to the semantic content that the model uses to understand the code structure and purpose. To test this, we conducted controlled experiments where we systematically removed these characters from the DocStrings and compared the model’s efficacy against a control group where these elements were retained.
Impact of Stop Words. We also investigated the role of stop words in DocStrings. In our study, we define the stop words are commonly considered to be filler words that do not carry much meaning, which are chosen by LLMLingua2 used in a 10% ratio we discussed in Section 3. Our experiments involved removing stop words from DocStrings to assess if the model’s efficacy would change. The comparison between the efficacy with and without stop words helped us understand the contribution of these words to the overall comprehension of different LLMs.
@@
Table 8. Efficacy Comparison whitin Different DocString Styles
DataSet Method DS-1.3B DS-6.7B CQ-7.3B CG-9.4B LA-8.0B Avg.
@@
Table 9. Efficacy comparison of Redundant Intruction
DataSet Method DS-1.3B DS-6.7B CQ-7.3B CG-9.4B LA-8.0B Avg.
@@
Findings and Implications. The findings from these experiments shed light on the importance of various elements within DocStrings. The removal of newline characters and tabs did not significantly impair the model’s efficacy, suggesting that these elements are more stylistic than functional in the context of code generation. The impact of stop words was also nuanced, with some instances showing a slight efficacy change, indicating that even seemingly inconsequential words might play a role in the model’s comprehension process. The removal of redundant instructions led to a modest improvement in efficacy, supporting our hypothesis that such phrases do not contribute positively to the model’s understanding and may even introduce noise.
@@ 6.3. Hyper-Parameter Analysis
Table 10. Efficacy Comparison of Different $\tau$ in CodeHarmony
DataSet $\tau$ Ratio DS-1.3B DS-6.7B CQ-7.3B CG-9.4B LA-8.0B Avg.
@@
Table 11. Efficacy Comparison of Different Base Models in CodeHarmony
DataSet Base Model DS-1.3B DS-6.7B CQ-7.3B CG-9.4B LA-8.0B Avg.
@@
(1) The hyperparameter $\tau$ significantly influences the performance of our method. As demonstrated in Table 10, different values of $\tau$ result in varying performance metrics across the CodeHarmony dataset. The optimal value of $\tau$ that maximizes performance may differ depending on the specific dataset and model architecture. For instance, a higher $\tau$ value like 0.999 seems to work better for the CodeQwen, CodeGeeX4, and Llama3.1 model, while a lower value like 0.985 could be more effective for the DeepSeekCoder-1.3B model. Given the good balance or optimal efficacy for the models and datasets, we set the hyperparameter $\tau$ to 0.999.
(2) The choice of base model also plays a critical role in the performance of our method. As shown in Table 11, different base models have different baseline performances. For the CodeHarmony dataset, the CodeGPT-py-adapted base model outperforms the other two models across all metrics, indicating that it is better suited for this specific dataset. This suggests that choosing the suitable base model is especially important for code generation tasks.
@@ 6.4. Threats to Validity
Threats to External Validity. The main external threat lies in the choice of datasets and models used in our study. To alleviate this, we selected six diverse datasets with high reputations to reflect the complexity of real-world scenarios. Importantly, our approach can be applicable to different programming languages. In terms of the choice of models, we selected six state-of-the-art models in code generation including DeepSeekCoder-1.3b, DeepSeekCoder-6.7b, CodeQwen1.5, CodeGeeX4, Llama3.1, and GPT-4o.
Threats to Construct Validity. Construct threats concern the performance metrics used to evaluate ShortenDoc and baselines. To evaluate the performance of models, we utilized Pass@1 metric which is commonly used in in the previous studies of code generation. Additionally, we used Compression Ratio, which is also widely used in similar studies, to measure the proportion of reduction in the length of the prompt after compression. Furthermore, we conducted a human study to analyze the qualitative impact of DocString compression.
@@ 7.2. Prompt Compression
Prompt Compression attempts to shorten the original prompts without losing essential information. Prompt Compression methods (Jiang et al., 2023b) can be grouped into three main categories : Token pruning and token merging, Soft prompt tuning methods and Information-entropy-based approaches. Token pruning and token merging need model fine-tuning or intermediate results during inference and have been used with BERT-scale models. For instance, Modarressi et al.(Modarressi et al., 2022) introduced AdapLeR which dynamically eliminates less contributing tokens through layers to achieve shorter lengths. Soft prompt tuning requires LLMs’ parameter fine-tuning to make them suitable for specific domains but not applicable to black-box LLMs. Wingate et al. (Wingate et al., 2022) used the framework of soft prompts to manipulate prompt compression. Mu et al. (Mu et al., 2024) proposed GIST, a manner to compresse arbitrary prompts into a smaller set of Transformer activations on top of virtual “gist” tokens. However, these methods are task-aware and usually tailored for specific tasks and compression ratios, which may limit their generalizability in real-world scenarios.
@@ Acknowledgements.
This work was partially supported by the National Natural Science Foundation of China (NSFC, No. 61972197 and No. 62372232), the Natural Science Foundation of Jiangsu Province (No. BK20201292), the Collaborative Innovation Center of Novel Software Technology and Industrialization, and the Short-term Visiting Program of Nanjing University of Aeronautics and Astronautics for Ph.D. Students Abroad (No. 240501DF16). T. Chen is partially supported by an oversea grant from the State Key Laboratory of Novel Software Technology, Nanjing University (KFKT2022A03).

### v3

@@ 5.3. RQ3: What is the impact of compressed prompts on developers? (Human Study)
RQ1 and RQ2 rely on automated evaluation metrics to compare the performance of different methods. In RQ3, we aim to gain more insights by conducting a human study to assess the impact of DocString compression.
To complement our analysis in RQ1 and RQ2, we first present a single example to provide deeper insights into how different compression methods affect code generation. While a single example cannot represent all scenarios, it helps illustrate the typical patterns we have observed across our experiments. Fig. 3 showcases a DocString from the Llama3.1 model within one representative example from the MBPP dataset, showing a DocString processed by different compression methods. We display the original DocString alongside its compressed versions, processed by four distinct methods. The original DocString led to semantically incorrect code generation by the Llama3.1 model. The Random compression method resulted in irrelevant code. Both Selective_Context and LLMLingua2 maintained the model’s output, yet failed to rectify the semantic inaccuracy. In stark contrast, our tool’s compression successfully steered the model towards semantically correct code generation.
To further evaluate the quality of the compressed DocStrings, we conducted a human study using three groups of DocStrings compressed by different methods (Selective_Context, LLMLingua2 and ShortenDoc).
@@
Figure 4. A Sample Questionnaire Used in Human Study. The ‘Prompt’ section shows the original uncompressed DocString, while the ‘Candidates’ section presents three compressed versions of the same DocString generated by different methods (SelectiveContext, LLMLingua2 and ShortenDoc). Evaluators were asked to rate each candidate on Informativeness (preservation of key information) and Comprehensibility (ease of understanding) using a 0-4 scale, without knowing which method produced each compression.
To conduct the human study, we recruited three evaluators with a strong background in Python programming (all are PhD candidates). We randomly selected 40 samples from each of the six datasets used in the previous experiments, resulting in a total of 240 samples for evaluation. For each sample, the evaluators assessed the original DocString along with three compressed versions generated by different methods.
To facilitate analysis and comparison across different data sources, we organized the samples into six groups according to their source datasets. Each group was evaluated by all three evaluators. The evaluators assessed each DocString anonymously based on two criteria: Informativeness and Comprehensibility. The evaluation scale ranged from 0 to 4 points, with higher scores indicating better quality. The final score for each criterion was calculated as the average of scores from all three evaluators.
The questionnaire used for the study is illustrated in Fig. 4. To maintain the quality of the human evaluation, the compressed DocStrings were presented in random order, ensuring that the evaluators were unaware of which method generated the DocString. Additionally, the evaluators were allowed to use Internet to look up unfamiliar concepts. To prevent fatigue and maintain concentration, we asked each evaluator to complete the assessment of one dataset (40 samples) over two days, with no more than 20 samples evaluated in any half-day session.
@@
Aspect Method Datasets
HumanEval CodeHarmony MBPP Subtle Creative BigCodeBench
@@
In terms of Informativeness, the ShortenDoc method consistently scores higher than the other two methods across most datasets. This suggests that ShortenDoc is adept at preserving the essential information within the DocStrings even after compression. In addition, the high Fleiss’ Kappa scores, ranging from 0.5 to 0.9, indicate a strong consensus among the evaluators regarding the Informativeness of the compressed DocStrings. This high agreement likely stems from the objective nature of this metric, where evaluators can clearly discern whether the compressed DocString retains the necessary information from the original.
In terms of Comprehensibility, ShortenDoc also shows superiority than the other two methods in Comprehensibility, indicating that the compressed DocStrings remain understandable to human readers. While the lower Fleiss’ Kappa scores for Comprehensibility, ranging from 0.1 to 0.3, suggest that there is less consistency among the evaluators in this metric compared to Informativeness.
The distinct scores in evaluator agreement between Informativeness and Comprehensibility can be attributed to their fundamental differences: (1) Subjectivity. Comprehensibility is inherently more subjective than Informativeness, as it depends on the evaluator’s personal ability to understand the compressed information. (2) Complexity of Docstrings. Some DocStrings, despite being compressed, may still contain complex information or jargon that is more challenging for some evaluators to grasp. (3) Diversity in Evaluator Background. The professional background and experience of the evaluators can influence their assessment of DocString comprehensibility. For instance, evaluators with specific domain knowledge might find certain terms or concepts more comprehensible.
@@ 6.1. The Impact of Method Name Quality
Table 7. Pass@1 comparison within Different Method Name Quality
DataSet Method LLMs
DS-1.3B DS-6.7B CQ-7.3B CG-9.4B LA-8.0B Avg.
@@
To empirically validate this hypothesis, we designed a controlled experiment. We systematically replaced high-quality method names with a generic placeholder, such as ‘foo’ (Yang et al., 2024b), across various datasets including HumanEval, CodeHarmony, MBPP, Subtle and Creative.
@@
The outcomes of this experiment are encapsulated in Table 7. The data clearly demonstrate that method names wield a significant influence on the Pass@1 score of DocString compression. Across the board, models exhibited a notable decline in Pass@1 score when high-quality method names were replaced with the generic ‘foo’. This finding suggests that a portion of the information compressed by ShortenDoc is redundant with the information already present in the method name.
Conversely, when low-quality method names were enhanced, we observed a consistent uptick in model Pass@1 score. This improvement underscores the notion that enriching the quality of method names not only aids in the comprehension of the code but also bolsters the robustness of ShortenDoc’s compression capabilities.
@@ 6.2. The Impact of DocString Style
Experiments with Newline and Tab Characters. In Section 4, during the preprocessing phase, we made an assumption that the removal of newline characters and tabs would not substantially affect the model’s code generation capabilities. This assumption was based on the idea that these characters, while useful for formatting, might not contribute to the semantic content that the model uses to understand the code structure and purpose. To test this, we conducted controlled experiments where we systematically removed these characters from the DocStrings and compared the model’s Pass@1 score against a control group where these elements were retained.
Impact of Stop Words. We also investigated the role of stop words in DocStrings. In our study, we define the stop words are commonly considered to be filler words that do not carry much meaning, which are chosen by LLMLingua2 used in a 10% ratio we discussed in Section 3. Our experiments involved removing stop words from DocStrings to assess if the model’s Pass@1 score would change. The comparison between the Pass@1 score with and without stop words helped us understand the contribution of these words to the overall comprehension of different LLMs.
@@
Table 8. Pass@1 Comparison whitin Different DocString Styles
DataSet Method LLMs
DS-1.3B DS-6.7B CQ-7.3B CG-9.4B LA-8.0B Avg.
@@
Table 9. Pass@1 comparison of Redundant Intruction
DataSet Method LLMs
DS-1.3B DS-6.7B CQ-7.3B CG-9.4B LA-8.0B Avg.
@@
[…]
The impact of stop words was also nuanced, with some instances showing a slight Pass@1 change, indicating that even seemingly inconsequential words might play a role in the model’s comprehension process. The removal of redundant instructions led to a modest improvement in Pass@1 score, supporting our hypothesis that such phrases do not contribute positively to the model’s understanding and may even introduce noise.
@@ 6.3. Hyper-Parameter Analysis
Table 10. Pass@1 Comparison of Different $\tau$ in CodeHarmony
DataSet $\tau$ Ratio LLMs
DS-1.3B DS-6.7B CQ-7.3B CG-9.4B LA-8.0B Avg.
@@
Table 11. Pass@1 Comparison of Different Base Models in CodeHarmony
DataSet Base Model LLMs
DS-1.3B DS-6.7B CQ-7.3B CG-9.4B LA-8.0B Avg.
@@
(1) The hyperparameter $\tau$ significantly influences the performance of our method. Since the hyperparameter $\tau$ is used to determine the threshold for token importance, it directly affects the number of tokens retained in the compressed DocString, i.e., the compression ratio. As demonstrated in Table 10, different values of $\tau$ result in varying performance metrics across the CodeHarmony dataset. The optimal value of $\tau$ that maximizes performance may differ depending on the specific dataset and model architecture. For instance, a higher $\tau$ value such as 0.999 seems to work better for the CodeQwen, CodeGeeX4 and Llama3.1 model, while a lower value like 0.985 could be more effective for the DeepSeekCoder-1.3B model. Given the good balance or optimal Pass@1 score for the models and datasets, we set the hyperparameter $\tau$ to 0.999.
A limitation in our hyperparameter analysis is that, while we conducted detailed $\tau$ experiments on CodeHarmony (as shown in Table 10), we did not perform exhaustive $\tau$ tuning across all datasets. The optimal $\tau$ value might vary across datasets due to their respective characteristics, and our choice of $\tau=0.999$ based on CodeHarmony may not be optimal for all datasets. Nevertheless, the experimental results demonstrate that $\tau=0.999$ consistently provides competitive Pass@1 scores across different datasets, as shown in Table 3. For instance, ShortenDoc achieves comparable or better Pass@1 than baseline methods in HumanEval (78.66%), CodeHarmony (63.40%) and MBPP (55.40%) when using the CodeQwen model.
We leave as future work a more comprehensive analysis of $\tau$ across different datasets to better understand its impact on compression performance and to develop dataset-specific tuning strategies.
(2) The choice of base model plays a critical role in the performance of our method. We find that different base models can lead to varying token importance rankings, indicating that the token sorting decisions are not entirely consistent across models. This observation raises interesting questions about the relationship between model characteristics and their ability to identify important tokens in DocStrings.
As shown in Table 11, different base models demonstrate varying baseline performances which can be attributed to their pre-training strategies: CodeGen-350M pre-trained specifically on code corpora outperforms GPT2 pre-trained solely on natural language corpora. Notably, CodeGPT-py-adapted achieves the best performance as it combines both approaches; it is built upon GPT2 and further pre-trained on code corpora. This suggests that the ability to understand both natural language and code context is crucial for effective DocString compression.
For the CodeHarmony dataset, while the CodeGPT-py-adapted base model outperforms the other models across all metrics, the token importance rankings produced by different models show notable variations. Despite these differences in token sorting, our method maintains robust performance across different base models, suggesting that the overall compression framework is resilient to variations in token importance assessment.
In summary, we advise users to set 0.999 as the default value of $\tau$ for balanced performance. If users wish to experiment with different compression ratios, they may gradually decrease this value. Meanwhile, we recommend using the CodeGPT-py-adapted model as the base model for DocString compression, as it consistently achieves the best performance across different datasets.
@@ 6.4. Threats to Validity
Threats to External Validity. The main external threat lies in the choice of datasets and models used in our study. For Python code generation, we selected six diverse datasets with high reputations to reflect the complexity of real-world scenarios. However, for cross-language evaluation, we relied on HumanEval-X, which consists of translations of the original HumanEval dataset. The reliance on translated versions of a single dataset, rather than naturally occurring code in different languages, might limit the generalizability of our findings. While this choice was constrained by the current availability of multilingual code generation benchmarks, future work could benefit from evaluating on native, language-specific datasets as they become available.
In terms of the choice of models, we prioritized open-source models for better reproducibility and community verification, selecting six representative models that span different scales (from 1.3B to 9.4B parameters) and architectures. These include specialized code models (DeepSeekCoder-1.3b, DeepSeekCoder-6.7b, CodeQwen1.5, CodeGeeX4) and general-purpose models (Llama3.1). While proprietary models such as GPT-4 might offer superior performance, their high API costs and frequent version updates could impact the stability and reproducibility of our results. Given the hardware constraint (single 3090 GPU) and the need for experimental consistency, we believe our current model selection provides meaningful insights into the effectiveness of our compression method across different scenarios.
Threats to Construct Validity. Construct threats concern the performance metrics used to evaluate ShortenDoc and baselines. To evaluate the performance of models, we utilized Pass@1 metric which is commonly used in in the previous studies of code generation. Additionally, we used Compression Ratio, which is also widely used in similar studies, to measure the proportion of reduction in the length of the prompt after compression.
Furthermore, we conducted a human study to analyze the qualitative impact of DocString compression, where the Informativeness and Comprehensibility metrics are used which might not capture all aspects of the DocString quality. We recruited evaluators with strong Python programming backgrounds and provided them with clear evaluation criteria. Additionally, we implemented measures to ensure evaluation quality, including anonymized compressed DocStrings, allowing evaluators to research unfamiliar concepts, and limiting their work load.
@@ 7.2. Prompt Compression
Prompt Compression attempts to shorten the original prompts without losing essential information. Prompt Compression methods (Jiang et al., 2023b) can be grouped into three main categories : Token pruning and token merging, Soft prompt tuning methods and Information-entropy-based approaches. Token pruning and token merging need model fine-tuning or intermediate results during inference and have been used with BERT-scale models. For instance, Modarressi et al.(Modarressi et al., 2022) introduced AdapLeR which dynamically eliminates less contributing tokens through layers to achieve shorter lengths. Soft prompt tuning requires LLMs’ parameter fine-tuning to make them suitable for specific domains but not applicable to black-box LLMs. Wingate et al. (Wingate et al., 2022) used the framework of soft prompts to manipulate prompt compression. Mu et al. (Mu et al., 2024) proposed GIST, a manner to compresse arbitrary prompts into a smaller set of Transformer activations on top of virtual “gist” tokens. Kobayashi et al. (Kobayashi et al., 2024) proposed the YORO paradigm, which aims to internalize database knowledge into model parameters to significantly reduce input length and simplify text-to-SQL input. However, these methods are task-aware and usually tailored for specific tasks and compression ratios, which may limit their generalizability in real-world scenarios.
@@ 8. Conclusion and Future Work
Furthermore, we aim to extend our research to general code completion tasks. Given the higher frequency of code completion in real-world software engineering compared to natural language to code generation, the motivation for prompt compression becomes even stronger in this context. This extension could potentially bring more significant practical benefits in terms of both computational and economic efficiency.
@@ Acknowledgements.
This work was partially supported by the National Natural Science Foundation of China (NSFC, No. 62372232), the Collaborative Innovation Center of Novel Software Technology and Industrialization, and the Short-term Visiting Program of Nanjing University of Aeronautics and Astronautics for Ph.D. Students Abroad (No. 240602DF16), and the A*STAR under its 2nd CSIRO and A*STAR: Research-Industry (2+2) Partnership Program (Award R24I5IR047). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the A*STAR. T. Chen is partially supported by an oversea grant from the State Key Laboratory of Novel Software Technology, Nanjing University (KFKT2023A04).